larry has an annual return of $213 from $3000 invested at simple interest, some at 5% the rest at 8%. how much was invested at each rate?
Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

 
Hi
Let x represent amount at 8%
.08x + .05(3000-x) = 213
x = 63/.03,amount at 8%
find (3000-x), amount at 5%

Answer by richwmiller(17219)

x+y=3000,
0.05*x+0.08*y=213
x=3000-y
0.05*(3000-y)+0.08*y=213
150-0.05y+0.08*y=213
0.03*y=63
y= 2100
x=3000-y
x=900
check
0.05*900+0.08*2100=213
45+168=213
213=213
ok